OpenStudy (anonymous):

The projected population of a certain ethnic group(in millions) can be approximated by p(t)=36.05(1.016)^t where t=0 corresponds to 2000 and 0<_t<_50. What is the instantaneous rate of change of the population when t=10?

OpenStudy (anonymous):

find d/dt and then set t =10

OpenStudy (anonymous):

can u break that down?

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@jennychan12

OpenStudy (jennychan12):

yeah what @tomo said for me, i would rewrite it as \[p(t) = 36.05(1.016)^t\] find d/dt Hint: \[\frac{ d }{ dx } (a^x) = a^x (\ln a )\]